Determine the force of gravitational attraction between the earth (m = 5.98 x 1024 kg) and a 63.0-kg man if the man is standing at sea level, a distance of 6.38 x 106 m from earth's center.


1
Expert's answer
2022-05-20T04:53:44-0400
"F=G\frac{m_1m_2}{r^2}"

"F=6.67*10^{-11}*\frac{5.98*10^{24}*63.0}{(6.38*10^6)^2}=617\:\rm N"